Jefferson Lab Weekly Briefs September 7, 2011
Jefferson Lab Weekly Briefs
September 7, 2011
12 GeV Upgrade
With 17 weeks into the 2011 Six Month Shutdown and 10 weeks remaining, 12 GeV installation work is progressing as planned. Installation activities will continue for approximately six weeks, and then the accelerator start-up and related commissioning activities will begin. Continued vigilance on safety remains a priority. In addition, 12 GeV logistical planning efforts are a continuous process to support fabrication and testing of components in preparation for installation during the next machine shutdown period.
Hall A has had a review of the status of the g2p installation. At this point, the completion date of the installation will mostly depend on the arrival time of equipment from vendors. The status report slides can be found online.
The large quadrupole magnet for the Hall C Moller polarimeter, which has instability problems, has been removed from the Hall C beamline by the mechanical installation group. An unused quadrupole, currently on the beamline, will be moved, replacing the functionality of the unstable magnet. Maintenance on the Q-weak high-power hydrogen target has been completed and the target has been moved back into the scattering chamber.
Much of last Monday was spent recovering systems from being shut down for the hurricane. The thermal relief valve for 2L24 was replaced and it was initially cooled down to 4 Kelvin then later to the 2K required for commissioning. The operator for the cryogenic valve downstream of 2L24 was replaced with a faster version, as the current design was too slow to be used with certain commissioning tests. Waveguide installation for 2L25 continues; pressure tests show no leaks. High-power amplifier testing of 2L25 continues to show good results. Magnet shim installations are complete, except for a few requiring adjustments. Refurbished dipole magnet polarity tests continue, with all polarity results as expected so far. The Fibercast repair of the South Access building cooling water flange continued. The South Linac service building cooling water system pressure tested okay. North Linac service building pipe inspection and installation of wear pads continues. Lastly, replacement has begun of the Machine Control Center (MCC) air conditioning compressor, which was damaged during the hurricane.
The FEL team stopped processing the gun at 320 kilovolts and heat cleaned the cathode over the weekend in preparation for running beam this week. The vault was reposted for running beam and radiofrequency was run in the cryomodules. The group intends to spend this week re-establishing beam in the injector. The ultraviolet wiggler passed its certification tests at STI Optronics and will be shipped back to JLab to arrive about Sept. 15.
Center for Theoretical and Computational Physics
In a new paper, the axial-vector current matrix elements between various mesons and baryons containing a single bottom quark were calculated in chiral perturbation theory. These results are necessary for extrapolations of lattice calculations of the corresponding matrix elements to physical values of the light quark masses. The calculations will also constrain predictions for the strong decay amplitudes for singly bottom baryons.
The installation group removed the Hall C Q-weak Moller quadruple and tested its coil. The results of this testing were that the coil failed and fixes are being examined. All of the dipoles in both the east arc and the west arc that were refurbished have been reinstalled. Additionally in the arcs, all beam pipes have been reconnected, put under vacuum, pumped down and leak checked. The cathode power supply in LN22 has been installed. In the west arc 10th pass, all XP four-meter dipoles and new QP-style quadruples that don't interfere with 6 GeV operations have been installed. All the warm region girders in the south linac have been installed except where the second C100 is scheduled to be installed during this down period.
The EES RF Group successfully ran RF power into waveguide shorts for the second new 12 GeV RF system on Sept. 2.
Environment, Safety, Health and Quality
This past Saturday, Sept. 3, marked the 20th anniversary of North Carolina’s worst industrial disaster: the Imperial Food Products plant fire in Hamlet that killed 25 workers and injured 56 others. After gas burners ignited a ruptured hydraulic line, panicked employees rushed for the emergency exit doors, only to find some of them padlocked. Rescue crews found bodies around the locked doors and in the freezer, where workers fled to escape the heat and smoke.
Computing and Networking Infrastructure
PC Purchase Program
Onsite TIAA-CREF Counseling Offered
Minority Undergraduate Research Assistantship
Dance Your Ph.D. Contest Call for Entries
Tracking Thomas at JLab
Congratulations to Harry Fanning, Marcy Stutzman and Doug Higinbotham, who were the first to correctly identify the Aug. 31 location. Honorable mentions go to David Anthony, Michelle Shinn, Ron Bartek and Mike Davis. Check out the Tracking Thomas webpage for a better view of his last location and this week's new mystery photo.
JLab Calendar of Events
Sept. 12: Safety Shoe vendor onsite